OpenStudy (anonymous):

Find all the real square roots of -9/16

Parth (parthkohli):

Observe...\[(-n)^2 = n^2\]\[n^2 = n^2\]A number has real square roots if and only if it is positive.

OpenStudy (ghazi):

\[\sqrt {\frac{ -9 }{ 16}}= \frac{ 3i }{ 4 }\]

Parth (parthkohli):

@ghazi: *real* square roots =)

OpenStudy (ghazi):

that doesnt exist for this rational number

OpenStudy (ghazi):

@ParthKohli

Parth (parthkohli):

Exactly.